Please help me solve this equation: 64x^2-128x-64
note: it is -128x not -28x

Answer by stanbon(75887)

solve this equation:
64x^2-128x-64 = 0
---
Factor:
64(x^2-2x-1) = 0
Divide by 64 to get:
x^2-2x-1 = 0
---------------------
Quadratic Formula:
x = [2 +- sqrt(4-4*1*-1)]/2
---
x = [2 +- sqrt(8)]/2
---
x = [2 +- 2sqrt(2)]/2
---
x = 1 + sqrt(2) or x = 1-sqrt(2)
=====================================
Cheers,
Stan H.